# XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

### v1

@@ 3 XPUTimer Design
Figure 3 illustrates the architecture of XPUTimer, which is deployed in Ant Group’s large-scale training cluster. XPUTimer is composed of two components: the tracing daemon and the diagnostic engine.
By automatically attaching a tracing daemon to each training process in LLM training jobs, XPUTimer enables a framework-agnostic and lightweight tracing mechanism. Specifically, the tracing daemon instruments only key code segments at the level of Python and C++ runtime, focusing on carefully selected Python APIs and GPU kernels. During instrumentation, events are injected to measure the latencies of these code segments. This approach draws on our experience hosting various large-scale LLM training jobs, targeting critical paths that commonly impact training efficiency.
@@
Then, the timing data collected by the daemons is transmitted to the diagnostic engine for holistic anomaly diagnostics. The engine employs a fast hang-error diagnostic method and leverages novel aggregated metrics to effectively identify slowdowns. The aggregated metrics encompass both macro-level metrics, such as training throughput, and micro-level metrics, including issue latency distribution of GPU kernels. Once the diagnostic engine identifies errors and slowdowns, the issues are routed to attribution teams, enabling swift resolution.
@@ 4 Lightweight Selective Tracing
To collect sufficient real-time data for anomaly diagnosis in the LLM training cluster, XPUTimer’s tracing daemon offers backbone-agnostic and lightweight full-spectrum tracing. Its design focuses on two key aspects: determining what information to collect and establishing how to collect it efficiently.
4.1 Key Segment Instrumentation
[…]
Figure 4: Instrumented key code segments in XPUTimer.
Therefore, XPUTimer selectively instruments code segments of key APIs and kernels to collect real-time information. This design is based on an insight into LLM training on large-scale GPUs: LLM training is predominantly dominated by a limited set of deep learning operators. These operators mainly include matrix multiplication and cross-GPU communication operators. Figure 4 presents the specific code segments instrumented by XPUTimer for efficient anomaly diagnostics.
As shown in the figure, the instrumented code segments can be broadly categorized into two groups. The first category involves intercepting key API calls, including those related to Python’s garbage collection (GC), PyTorch’s dataloader, and GPU synchronization. These APIs are carefully selected based on empirical insights into performance issues and optimization opportunities. These insights are detailed in §5.2, with corresponding cases discussed in §7.
The second category focuses on intercepting critical GPU computation and communication kernels executed at the C++ runtime level. These kernels, primarily provided by optimized libraries[4, 6, 46, 5], account for the majority of the workload during large-scale training. Additionally, there are customized kernels developed by the infrastructure team.
Notably, the above design enables XPUTimer to support backbone-agnostic and extensible tracing capabilities. Extending tracing capabilities for Python-related APIs is straightforward, requiring only the configuration of the specific environment variables in the training scripts, as shown below. {minted}[ frame=none, obeytabs=true, framesep=0mm, baselinestretch=0.8, fontsize=, xleftmargin=1.6em, breaklines, escapeinside=||, ]shell export TRACED_PYTHON_API="torch.cuda@synchronize" Meanwhile, intercepting C++ kernels necessitates explicit registration through a C++ interface. This requirement is feasible, as the infrastructure team takes charge of the development of both these customized operators and XPUTimer, ensuring seamless integration and functionality.
4.2 Timing in the Background
@@
For synchronous Python API calls, FLARE directly records their start and end timestamps and forwards them to the timing manager. For GPU kernels, which execute asynchronously, FLARE injects CUDA events[47] after an interception to record execution status. These events are enqueued for further processing. The timing manager queries the status of the queued events in the background, avoiding any disruption to the training thread. Additionally, during GPU kernel interception, FLARE extracts input specifications, such as memory layout, to support subsequent anomaly diagnostics in §5.2.1.
Figure 5: Intercepting and timing the training in the background.
As training progresses, the timing manager proactively transmits all real-time data to XPUTimer’s diagnosis engine. By employing key segment instrumentation and running timing tasks in the background, XPUTimer minimizes both computing resource and memory overhead, ensuring efficient data collection for real-time anomaly diagnostics. A detailed evaluation of XPUTimer’s real-time overhead is provided in §6.1.
@@ 5 Anomaly Detection and Diagnosis
Using the real-time data collected by the tracing daemon, XPUTimer’s diagnostic engine identifies and analyzes anomalies encountered during distributed LLM training. In this section, we present XPUTimer’s holistic diagnostic workflow for addressing two common anomaly symptoms: errors and slowdowns.
@@ 5.1 Fast Runtime Error Diagnosis
As shown in the left of Table 1, errors encountered at the beginning of a training job are typically caused by bugs in the training scripts, which can often be addressed easily by the algorithm teams and infrastructure team. However, diagnosing errors that occur during the training progress is more challenging and critical. Such errors often stem from issues like operating system crashes, GPU failures, or network disruptions, which can generally be resolved by isolating the problematic machines and restarting the training job.
A typical symptom associated with these errors is the hanging of the training job. Training LLMs across numerous GPUs in a distributed manner inherently relies on the coordination of training processes. When the aforementioned errors occur, they rarely affect all training processes simultaneously. In this context, XPUTimer focuses on rapidly diagnosing hang errors by identifying faulty machines. Then, XPUTimer routes this information to the operations team, enabling the training job to restart with healthy machines.
Specifically, XPUTimer’s diagnostic engine first detects hang errors by examining the status of tracing daemons. The tracing daemon operates in the background of the training thread and continuously queries events recorded during job execution. If it fails to confirm the completion of an event within a predefined timeout interval, it proactively reports a potential hang error to the diagnostic engine. Similarly, if a tracing daemon does not transmit any real-time data within the specified timeout interval, the diagnostic engine also interprets this as an indication of a hang error.
After hang errors are reported, they are classified as either communication or non-communication errors. XPUTimer diagnoses these errors in two steps: first, a coarse-grained diagnosis through call stack analysis; and second, a fine-grained diagnosis using intra-kernel tracing
@@
This diagnosis is used to identify problematic machines encountering non-communication errors. Figure 6 illustrates an example of hang-error diagnosis via call stack analysis. As shown in the left of Figure 6, when the training process of rank-00 crashes or is suspended due to a non-communication error, it halts at a call stack corresponding to a non-communication function. In contrast, the training processes of other ranks continue executing correctly and eventually stop at a call stack associated with a communication-related function that depends on coordination with rank-00. In this scenario, the machine associated with rank-00 is identified as the source of the error. It should be noted that, although these non-communication errors may cause direct crashes, the call stack analysis could also locate the faulty machine.
However, communication hang errors cannot be identified through call stack analysis. As shown in the right of Figure 6, the training processes of all ranks terminate at the same call stack corresponding to a communication function, such as allreduce or allgather. In this scenario, there are no distinct differences between ranks based on call stack analysis.
@@
We further investigate the symptoms of communication hang errors and obtain two observations. Firstly, some communication hang errors generate error logs. For instance, if the link between RDMA NICs breaks, an error code of 12 is produced. Secondly, more hang errors result in an endless loop within the launched communication kernels, ultimately leading to job termination after a predefined timeout. To identify the unhealthy machine responsible for such errors, a straightforward solution is to perform a binary search by executing communication tests across all involved GPUs. This approach has a complexity of $O(\log N)$ and requires hours to pinpoint the faulty machine among thousands of GPUs [14].
Diagnosis using intra-kernel tracing.
Faced with this problem, XPUTimer introduces a minute-level diagnostic approach using intra-kernel tracing. This intra-kernel tracing leverages CUDA-GDB, the debugging tool for CUDA programming.
Specifically, XPUTimer’s diagnostic engine instructs the tracing daemon to attach the halted training processes with CUDA-GDB before terminating them. Once attached, the tracing daemon executes a script capable of automatically extracting detailed communication statuses to identify unhealthy machines. Figure 7 depicts an example of diagnosing communication hang errors in a hanging ring-allreduce kernel.
In the ring-allreduce kernel, each thread block is responsible for transmitting data between linked adjacent ranks within the kernel’s constructed ring. The data are split into chunks and thread blocks of adjacent ranks work together to transmit the chunks step by step. Thus, XPUTimer could retrieve the register values corresponding to the loop steps used for data transmission between linked ranks. Theoretically, the connection with the minimum step reveals the related GPUs experiencing errors. This intra-kernel tracing process is performed in parallel across all involved GPUs. As a result, its complexity is $O(1)$, enabling completion within a few minutes.
XPUTimer then routes the diagnostic information for detected errors to the operations team, assisting with tasks such as isolating faulty machines and restarting the training job.
@@ 5.2 Aggregation for Slowdown Diagnosis
As demonstrated in §2.2, slowdowns can be attributed to changes across the entire training stack. Meantime, slowdowns caused by software changes introduced by the algorithm and infrastructure teams are often subtle and challenging to detect. Identifying these changes typically requires comparisons across historical training jobs and prior training steps. In contrast, hardware changes, such as GPU underclocking or network jitter, are more apparent and can be detected solely through comparisons across training steps.
@@
To holistically identify these anomalies, XPUTimer aggregates real-time data collected from the tracing daemon into five primary metrics, shown in Figure 8. These metrics are based on the consensus that a “healthy” training pipeline should exhibit a timeline saturated with GPU kernels dedicated to either computation or communication. Computation kernels should achieve high FLOPS, while communication kernels are expected to utilize high bandwidth. Any deviations from these characteristics point to idle GPU resources, signaling potential slowdowns in training jobs. Of the five metrics, three are commonly used in existing works [2, 14], while the other two are newly introduced by XPUTimer.
Figure 7: An example of diagnosing communication hang errors in a ring-allreduce kernel using intra-kernel tracing.
Figure 8: A timeline of a distributed training job annotated with aggregated metrics used for diagnosing slowdowns in XPUTimer.
5.2.1 Diagnosing Obvious Slowdown
1 Training throughput for detecting slowdown.
Training throughput is the most straightforward metric for detecting slowdowns. XPUTimer measures training throughput by timing the rate at which input data is consumed by the training pipeline. This is achieved by instrumenting the dataloader API of Pytorch. As a macro performance metric, training throughput directly reflects slowdowns in training efficiency through comparison to historical training jobs and between training steps of the same job. However, XPUTimer cannot diagnose the specific factors contributing to the slowdown. To address this, XPUTimer relies on the following four micro metrics to further investigate the underlying causes.
2 FLOPS for slow critical kernel.
XPUTimer monitors the FLOPS of instrumented critical computation kernels, leveraging timing data and input layout. By comparing the FLOPS of identical kernels across different ranks, XPUTimer diagnoses GPUs that exhibit poor computational performance, often caused by issues like GPU underclocking. Machines affected by GPU underclocking are then routed to the operations team for isolation. Additionally, by analyzing FLOPS, XPUTimer identifies un-optimized kernels in training jobs, particularly those with large input sizes but low FLOPS. These anomalies are detected without interrupting training jobs and are subsequently routed to the infrastructure team for further investigation. Notably, when analyzing FLOPS data, XPUTimer accounts for the impact of communication kernels that overlap with computation kernels.This ensures that computation kernels with falsely low FLOPS are not mistakenly flagged.
3 Bandwidth for slow connection.
XPUTimer monitors the bandwidth of communication kernels. A communication operator requires launching the communication kernels on all ranks. Since variations in kernel-issue timestamps exist across different ranks, XPUTimer calculates the communication bandwidth by utilizing the start and end timestamps of the final communication kernels issued across all participating ranks. The captured communication bandwidth is compared with offline profiled data. If low-bandwidth communication is detected, XPUTimer conducts a communication test using binary search to pinpoint machines experiencing issues such as network congestion. These slowdowns are then identified and routed to the operations team for resolution.
5.2.2 Diagnosing Obscured Slowdown
While the above three metrics ensure that both critical computation and communication GPU kernels operate at high performance, they do not cover the less critical operations, such as various CPU operations and element-wise activation GPU kernels. Meantime, XPUTimer’s selective key segment instrumentation also omits the monitoring of these operations.
@@ 5.2.2 Diagnosing Obscured Slowdown
To diagnose their potential contributions to slowdowns, we further classify these not-instrumented operations into three categories: intra-step CPU operations, inter-step CPU operations, and minority GPU kernels. Intra-step CPU operations and inter-step CPU operations differ due to their occurrences within the timeline of training steps. Minority GPU kernels refer to those GPU kernels that often occupy little GPU computation resources. Specifically, two metrics are introduced for the diagnostics: issue latency distribution for intra-step CPU operations and void percentage for inter-step CPU operations and minority GPU kernels.
Figure 9: Runtime overhead in terms of latency with various models, backbones, and number of GPUs.
@@
In a well-optimized parallel backbone, only the necessary intra-step CPU operations for launching GPU kernels or coordinating the training processes are expected. However, algorithm teams may inadvertently introduce unnecessary GPU synchronizations when modifying the LLM model. Meantime, certain function calls, such as GC [2, 9], may be implicitly triggered by the Python runtime. These intra-step CPU operations can occur repeatedly during the model’s forward pass, bringing considerable overhead. In such cases, these operations cause an anomaly known as a kernel-issue stall, leading to GPU idle time within the training step.
@@
–1 in Figure 8 shows the example of Python runtime GC. In the figure, the Python runtime GC stalls the CPU thread and causes the lagging of GPU kernels on rank-1. Although the communication kernel on rank-0 is issued without stalling, it simply waits for the one on rank-1, ultimately causing the overall training speed to decline. 4–2 in Figure 8 shows an example of unnecessary GPU synchronization introduced by the developers from the algorithm teams. As all ranks wait for the completion of communication kernels, the kernel issue of follow-up kernels is stalled and not overlapped with GPU computation. When such unnecessary synchronization occurs repeatedly across the model’s forward pass, it ultimately results in a slowdown of the training speed.
Originally, detecting these anomalies of kernel-issue stall requires investigating the aggregated timeline with much human effort. Faced with this issue, XPUTimer proposes a new metric, named issue latency distribution, for diagnosing this issue without human intervention. Kernel-issue latency is defined as the time elapsed between the kernel’s issue timestamp and the start timestamp of its execution on the GPU. Based on our observation of anomalies of kernel-issue stall, the kernel-issue latencies of unhealthy training jobs should be much shorter than those of a healthy training job.
By monitoring runtime issue latency distribution, and comparing it with the historical data, XPUTimer could identify the anomaly of kernel-issue stall. Then, XPUTimer routes them to algorithm and infrastructure teams for resolution, as they are commonly software issues. §6.3 and §7.2 demonstrate the effectiveness of issue latency distribution in diagnosing kernel-issue stalls, even when the slowdown is minimal.
@@
While the tracing daemon only instruments the critical operators, inter-step CPU operations and minority GPU kernels both manifest as empty time slots in the visualized timeline, as shown in Figure 8. Consequently, XPUTimer introduces a metric, termed the void percentage, to identify slowdowns caused by these factors.
As for inter-step CPU operations, as depicted by 5–2 in Figure 8, XPUTimer measures the latency between the last kernel preceding the dataloader and the first kernel following the same dataloader. XPUTimer then computes the void percentage for inter-step CPU operations using the following equation:
@@
As for minority GPU kernels, as shown by 5–1 in Figure 8, XPUTimer first automatically detects empty slots where GPU kernels are launched but remain un-executed. These empty slots signify that the GPUs are occupied by kernels outside the scope of XPUTimer’s tracing mechanism. XPUTimer subsequently accumulates these slots for each training step and computes the void percentage using the following equation:
@@
When the void percentages ($V_{inter}$ and $V_{minority}$) surpass the predefined thresholds for a specific parallel backbone, XPUTimer annotates the training job with potential slowdowns attributed to inter-step CPU operations or minority GPU kernels. Then, XPUTimer notifies the algorithm and infrastructure team for further investigation.

### v2

@@ 3 Flare Design
In this paper, we propose Flare, a diagnostic framework for diverse large-scale LLM training. It places particular emphasis on diagnosing performance regression automatically from algorithm and infrastructure teams. Figure 2 illustrates the architecture of Flare, deployed in Ant Group’s large-scale training cluster. Flare is composed of two components: the tracing daemon and the diagnostic engine.
@@
By automatically attaching a tracing daemon to each training process in LLM jobs, Flare provides a lightweight tracing mechanism. The daemon is backend-extensible but also enables the collection of sufficient runtime data across the entire training stack. For each job, runtime data such as function timing events and call stacks are intercepted for monitoring and diagnostics. The diagnostic engine employs a fast hang-error diagnostic method and leverages novel aggregated metrics to effectively identify slowdowns, especially regressions. The aggregated metrics encompass both macro-level metrics, such as training throughput, and micro-level metrics, including issue latency distribution of GPU kernels.
The overall diagnostic pipeline is as follows:
①Fast hang-error diagnosis enables error detection and analysis. Fail-slows are detected through macro metrics and validated by micro metrics. Errors and fail-slows are then routed to the operations team for independent handling.
②Regressions are detected by monitoring the proposed micro metrics and diagnosed via Python API analysis. Flare routes potential regressions with narrowed root causes to the algorithm or infrastructure team.
③The algorithm, infrastructure, and operations teams collaborate only when anomalies cannot be resolved independently by the routed team.
@@ 4 Lightweight Selective Tracing
The runtime data collection overhead primarily stems from high memory usage rather than interference with computing resources, since profiling APIs like CUPTI[cupti2024] can operate in a background thread. For instance, profiling a Llama-70B model trained on 512 H800 GPUs using PyTorch’s built-in profiler produces a log file of $5.5\text{GB}$ (in JSON format, compressed to $451\text{MB}$) for each training step. This substantial memory overhead renders such arbitrary profiling methods impractical for continuously collecting real-time data.
Therefore, Flare selectively instruments code segments of key APIs and kernels to collect real-time information. This design is based on an insight into LLM training on large-scale GPUs: LLM training is predominantly dominated by a limited set of deep learning operators. These operators mainly include matrix multiplication and cross-GPU communication operators. In the rest of this section, we introduce the way Flare achieves full-stack and backend-extensible tracing through employing two techniques: plug-and-play instrumentation and timing with stack reconstruction.
4.1 Plug-and-Play Instrumentation
Figure 3: Instrumented key code segments in Flare.
Full-stack, backend-extensible tracing requires plug-and-play instrumentation that captures the required runtime data without intrusive backend changes. Figure 3 shows the code segments instrumented by Flare, which are widely used in backend codebases and broadly fall into two groups. The first intercepts key API calls, including Python’s garbage collection (GC), PyTorch’s dataloader, and GPU synchronization. The second targets critical GPU computation and communication kernels executed at the C++ runtime level. These kernels dominate the workload in large-scale training.
Intercepting C++ functions is straightforward with backend extensibility, as mechanisms like LD_PRELOAD can hook any dynamically linked C++ functions. In contrast, intercepting Python APIs is more challenging when enforcing backend extensibility. A common method for tracing is to provide interfaces such as Python decorators to wrap APIs in backend. E.g., MegaScale supports FSDP by patching PyTorch, which requires intruding into each backend for full-stack support.
To avoid such intrusiveness for backend extensibility, Flare adopts a CPython-based intercepting mechanism. It maintains a list of tracing-required APIs for each backend. When a training job starts, Flare imports these APIs and retrieves their bytecode. During long training runs, it intercepts them directly using CPython’s profiling API PyEval_SetProfile based on the bytecode.
The above design allows Flare to instrument either Python or C++ functions in a plug-and-play manner. For easy plug-in, Flare does not modify any backend codebase before tracing. For easy play, Flare exposes user-friendly interfaces for tracing new functions. Specifically, users only need to configure environment variables before launching jobs, such as export TRACED_PYTHON_API=‘‘torch.cuda@synchronize’’.
Intercepting C++ functions requires explicit registration through a C++ interface. This is practical because the infrastructure team develops both these C++ functions and Flare, ensuring seamless integration and functionality.
4.2 Timing with Stack Reconstruction
@@
For synchronous Python API calls, FLARE directly records their start and end timestamps and forwards them to the timing manager. For GPU kernels, which execute asynchronously, FLARE injects CUDA events[cudaevents2024] after an interception to record execution status. These events are enqueued for further processing. The timing manager queries the status of the queued events in the background, avoiding any disruption to the training thread. Additionally, during GPU kernel interception, FLARE extracts input specifications, such as memory layout, to support subsequent anomaly diagnostics.
As training progresses, the timing manager proactively streams real-time data to Flare’s diagnostic engine. However, since plug-and-play instrumentation uses separate methods for Python APIs and C++ functions, it omits the call stack linking them, which is essential for anomaly diagnostics. Fortunately, we record not only their durations but also their start and end timestamps. Using this information, the tracing thread reconstructs the call-stack relationships before sending the data to the diagnostic engine.
Figure 4: Timing the training in the background.
@@ 5 Anomaly Detection and Diagnosis
In this section, we present Flare’s automated diagnosis of anomalies: runtime errors, fail-slows, and performance regressions. Among these, Flare pays particular attention to regressions, which have not been explored in prior work.
@@ 5.1 Fast Runtime Error Diagnosis
Runtime errors often stem from issues like operating system crashes, GPU failures, or network disruptions, which can generally be resolved by isolating the problematic machines and restarting the training job. A typical symptom associated with these errors is the hanging of the training job. Training LLMs across numerous GPUs in a distributed manner inherently relies on the coordination of training processes. When the aforementioned errors occur, they rarely affect all training processes simultaneously. The training process hangs if a GPU fails. In this context, Flare focuses on rapidly diagnosing hang errors by identifying faulty machines. Then, Flare routes this information to the operations team, enabling the training job to restart with healthy machines.
Specifically, the diagnostic engine detects hang errors by examining the status of tracing daemons. The daemon operates in the background of the training thread and continuously queries events recorded during job execution. If it fails to confirm the completion of an event within a predefined timeout interval, it proactively reports a potential hang error to the diagnostic engine. If a tracing daemon transmits none real-time data within the specified timeout interval, the diagnostic engine also interprets this as an indication of a hang error.
After hang errors are reported, they are classified as either communication or non-communication errors. Flare diagnoses these errors in two steps: first, a coarse-grained diagnosis through call stack analysis; and second, a fine-grained diagnosis using intra-kernel inspecting.
@@
This diagnosis is used to identify problematic machines encountering non-communication errors. Figure 5 illustrates an example of hang-error diagnosis via call stack analysis. As shown in the left of Figure 5, when the training process of rank-0 crashes or is suspended due to a non-communication error, it halts at a call stack corresponding to a non-communication function. In contrast, the training processes of other ranks continue executing correctly and eventually stop at a call stack associated with a communication-related function that depends on coordination with rank-0. In this scenario, the machine associated with rank-0 is identified as the source of the error.
However, communication hang cannot be identified through call stack analysis. As shown in the right of Figure 5, the training processes of all ranks terminate at the same call stack of a communication function, such as allreduce or allgather. We further investigate the symptoms of communication hang errors and obtain two observations. Firstly, some communication hang errors generate error logs. For instance, if the link between RDMA NICs breaks, an error code of 12 is produced. Secondly, more hang errors result in an endless loop within the launched communication kernels, ultimately leading to job termination after a predefined timeout.
@@
To identify the faulty machine responsible for such errors, a common approach is to terminate the training process and run NCCL tests across all GPUs. However, in large-scale training where multiple model parallelisms–such as tensor [shoeybiMegatronLMTraining], pipeline [narayananPipeDreamGeneralized], and expert parallelism [lepikhinGshardScaling]–are combined, the NCCL tests must span all configured communication groups. In our experience, this exhaustive and blind search can take over half an hour at the thousand-GPU scale.
Figure 6: Diagnosing hang errors via intra-kernel inspecting.
Figure 7: A timeline of a distributed training job annotated with aggregated metrics used for diagnosing slowdowns in Flare.
Diagnosis using intra-kernel inspecting.
Faced with this problem, Flare introduces a minute-level diagnostic approach using intra-kernel inspecting with $\mathcal{O}(1)$ complexity. Instead of stopping the hanged training processes and blindly probing with NCCL tests, Flare fully utilizes real-time data at the error point for diagnosis. Specifically, the intra-kernel inspecting leverages CUDA-GDB, the debugging tool for CUDA programming. It instructs the tracing daemon to attach to the halted training processes using CUDA-GDB. Once attached, the tracing daemon executes a script capable of automatically extracting detailed communication statuses to identify unhealthy machines.
Figure 6 depicts an example of diagnosing communication hang errors in a hanging ring-allreduce kernel. In the ring-allreduce kernel, each thread block is responsible for transmitting data between linked adjacent ranks within the kernel’s constructed ring. The data are split into chunks and thread blocks of adjacent ranks work together to transmit the chunks step by step. Thus, Flare retrieves the register values corresponding to the loop steps used for data transmission between linked ranks. Theoretically, the connection with the minimum step reveals the related GPUs experiencing errors. This intra-kernel inspecting process is performed in parallel across all involved GPUs. As a result, its complexity is $O(1)$, enabling completion within a few minutes.
Flare then routes the diagnostic information for detected errors to the operations team, assisting with tasks such as isolating faulty machines and restarting the training job. Notably, this mechanism does not degrade the performance of the training job. The tracing daemon directly inspects the SASS code to retrieve register values. Thus, Flare does not require recompiling the NCCL library with debugging information enabled, which would disables several compiler optimizations. Moreover, Flare attaches to the training processes only after the job hangs, without interfering with normal execution.
@@ 5.2 Aggregation for Slowdown Diagnosis
To holistically identify fail-slow and regression slowdowns, Flare aggregates real-time data from the tracing daemon into five primary metrics, shown in Figure 7. These metrics reflect the consensus that a “health” training pipeline should exhibit a timeline saturated with GPU kernels dedicated to computation or communication. Computation kernels should deliver high FLOPS, while communication kernels should sustain high bandwidth. Deviations from these characteristics indicate idle GPU resources, signaling potential slowdowns in training jobs. In the rest of this section, we introduce the five metrics and explain how they are used to detect and narrow down the root causes of fail-slows and regressions.
5.2.1 Detecting Fail-slows with Macro Metric
1 Training throughput for detecting fail-slows.
Fail-slows caused by low-level hardware changes, such as GPU underclocking or network jitter, are apparent and can be detected solely through comparisons across training steps. To this end, Flare measures training throughput by timing the rate at which input data is consumed by the training pipeline. It is achieved by instrumenting the dataloader API of Pytorch.
5.2.2 Detecting Regression with Micro Metrics
Regressions caused by software changes introduced by algorithm and infrastructure teams are often subtle and difficult to detect. Identifying them usually requires manual comparisons across historical training jobs. However, directly comparing training throughput between monitored and historical jobs does not reliably reveal regressions, since model size, input datasets, and other factors may differ significantly. To address this, Flare detects regressions using micro metrics. This is because when macro metrics do not show sudden changes, abnormalities in micro metrics often indicate potential persistent performance regressions.
2 FLOPS and 3 bandwidth.
FLOPS and bandwidth capture the performance of instrumented computation and communication kernels, and can be used to detect regressions. For example, computation kernels with large input sizes but much lower FLOPS than the theoretical value indicate a regression in computation performance.
Flare monitors the FLOPS of instrumented critical computation kernels, leveraging timing data and input layout. It also monitors the bandwidth of communication kernels. A communication operator requires launching the communication kernels on all ranks. Since variations in kernel-issue timestamps exist across different ranks, Flare calculates the communication bandwidth by utilizing the start and end timestamps of the final communication kernels issued across all participating ranks.
Notably, when analyzing FLOPS and bandwidth to detect regressions, Flare accounts for the overlap of communication and computation kernels, which is common in training MoE-based LLMs [deepseek-aiDeepSeekV3Technical]. This ensures that computation kernels with falsely low FLOPS are not mistakenly flagged.
@@ 5.2.2 Detecting Regression with Micro Metrics
While FLOPS and bandwidth ensure that both critical computation and communication GPU kernels operate at high performance, they do not cover the less critical operations, such as various CPU operations and element-wise activation GPU kernels. Meantime, Flare’s tracing also omits the monitoring of these operations. To detect their potential contributions to regressions, we further classify these not-instrumented operations into three categories: intra-step CPU operations, inter-step CPU operations, and minority GPU kernels. Intra-step CPU operations and inter-step CPU operations differ due to their occurrences within the timeline of training steps. Minority GPU kernels refer to those GPU kernels that often occupy little GPU computation resources.
Specifically, two metrics are introduced for regression detection: issue latency distribution for intra-step CPU operations and void percentage for inter-step CPU operations and minority GPU kernels.
@@
In a well-optimized parallel backend, only the necessary intra-step CPU operations for launching GPU kernels or coordinating the training processes are expected. However, algorithm teams may inadvertently introduce unnecessary GPU synchronizations when modifying the LLM model. Meantime, certain function calls, such as GC [jiangMegaScaleScaling, shoeybiMegatronLMTraining], may be implicitly triggered by the Python runtime. These intra-step CPU operations can occur repeatedly during the model’s forward pass, bringing considerable overhead. In such cases, these operations cause a regression anomaly known as a kernel-issue stall, leading to GPU idle time within the training step.
@@
–1 in Figure 7 shows the example of Python runtime GC. In the figure, the Python runtime GC stalls the CPU thread and causes the lagging of GPU kernels on rank-1. Although the communication kernel on rank-0 is issued without stalling, it simply waits for the one on rank-1, ultimately causing the overall training speed to decline. 4–2 in Figure 7 shows an example of unnecessary GPU synchronization introduced by the developers from the algorithm teams. As all ranks wait for the completion of communication kernels, the kernel issue of follow-up kernels is stalled and not overlapped with GPU computation. When such unnecessary synchronization occurs repeatedly across the model’s forward pass, it ultimately results in a regression of the training speed.
Originally, detecting these anomalies of kernel-issue stall requires investigating the aggregated timeline with much human effort. Faced with this issue, Flare proposes a new metric, named issue latency distribution, for diagnosing this issue without human intervention. Kernel-issue latency is defined as the time elapsed between the kernel’s issue timestamp and the start timestamp of its execution on the GPU. Based on our observation of regression due to kernel-issue stall, the kernel-issue latencies of unhealthy training jobs should be much shorter than those of a healthy training job.
Ahead of deployment, Flare learns healthy kernel issue distributions from historical data for each backend type and cluster scale. Flare use the maximum Wasserstein distance [ramdas2017wasserstein] between these healthy distributions as a threshold. At runtime, Flare compares the collected data against the healthy distribution by computing the Wasserstein distance between them. A warning is triggered when the distance exceeds the learned threshold.
@@
While the tracing daemon only instruments the critical operators, inter-step CPU operations and minority GPU kernels both manifest as empty time slots in the visualized timeline, as shown in Figure 7. Consequently, Flare introduces a metric, termed the void percentage, to identify slowdowns caused by these factors.
As for inter-step CPU operations, as depicted by 5–2 in Figure 7, Flare measures the latency between the last kernel preceding the dataloader and the first kernel following the same dataloader. Flare then computes the void percentage for inter-step CPU operations using the following equation:
@@
As for minority GPU kernels, as shown by 5–1 in Figure 7, Flare first automatically detects empty slots where GPU kernels are launched but remain un-executed. These empty slots signify that the GPUs are occupied by kernels outside the scope of Flare’s tracing mechanism. Flare subsequently accumulates these slots for each training step and computes the void percentage using the following equation:
@@
When the void percentages ($V_{inter}$ and $V_{minority}$) surpass the predefined thresholds for a specific parallel backend, Flare annotates the training job with potential regressions attributed to inter-step CPU operations or minority GPU kernels.
5.2.3 Diagnosing Root Causes for Fail-slows
While changes in training throughput between steps of the same job indicate fail-slows, Flare cannot directly diagnose the specific factors behind these slowdowns. Following prior work, Flare investigates the root causes using two micro metrics mentioned above: FLOPS and bandwidth.
By comparing the FLOPS of identical kernels across different ranks, Flare diagnoses GPUs that exhibit poor computational performance, often caused by issues like GPU underclocking. Machines affected by GPU underclocking are then routed to the operations team for isolation. The captured communication bandwidth is compared with offline profiled data. If low-bandwidth communication is detected, Flare conducts a communication test using binary search to pinpoint machines experiencing issues such as network congestion. These slowdowns are then identified and routed to the operations team for resolution.
5.2.4 Diagnosing Root Causes for Regressions
Pinpointing the root causes of regressions in computation and communication kernels is straightforward. Flare directly forwards the traced data, such as input shape and layout, to the infrastructure team for resolution. When Flare detects regressions related to kernel-issue stalls, it further narrows down the root causes by analyzing the invocation of relevant Python APIs recorded by the tracing daemon. Specifically, Flare checks for APIs such as Python GC invoked just before communication kernels with abnormal issue distributions. The same approach applies to regressions detected by void percentage. If relevant APIs are found, Flare directly pinpoints the potential root causes and first routes them to the corresponding algorithm teams. If the algorithm teams cannot resolve them independently, or if no relevant Python APIs are found, Flare forwards the regressions to the infrastructure team for further investigation. In this case, both algorithm and infrastructure teams validates and resolves regressions with the root causes already narrowed down.